Oriole Corp. purchased machinery for $381,000 on May 1, 2020. It is estimated that it will have a useful life of 10 years, salvage value of $18,000, production of 290,400 units, and working hours of 25,000. During 2021, Oriole Corp. uses the machinery for 2,650 hours, and the machinery produces 30,600 units.

From the information given, compute the depreciation charge for 2021 under each of the following methods. (Round intermediate calculations to 2 decimal places, e.g. 5.25 and final answers to 0 decimal places, e.g. 45,892.)

(a)

Straight-line

$enter a dollar amount

(b)

Units-of-output

$enter a dollar amount

(c)

Working hours

$enter a dollar amount

(d)

Sum-of-the-years'-digits

$enter a dollar amount

(e)

Declining-balance (use 20% as the annual rate)

$enter a dollar amount
Answer #1

Solution

Oriole Corp

Computation of Depreciation charge for 2021under each of the following methods:

a

Straight line

$36,300

b

Units-of-Output

$38,250

c

Working hours

$38,478

d

Sum-of-the-years’-digits

$59,400

e

Declining Balance

$66,040

Computations:

  1. Straight line –

Depreciation expense = depreciable base x 1/useful life

Depreciable base = cost – salvage value

Cost = $381,000

Salvage value = $18,000

Depreciable base = 381,000 – 18,000 = $363,000

Useful life = 10 years

Depreciation expense = 363,000/10 = $36,300

Straight line depreciation charge for 2021 = $36,300

  1. Units-of-Output –

Depreciation expense = depreciable rate x number of units

Depreciable rate = depreciable base/estimated production

Depreciable base = cost – salvage value

Cost = $381,000

Salvage value = 18,000

Depreciable base = 381,000 – 18,000 = $363,000

Production = 290,400 units

Depreciable rate per unit = $363,000/290,400 = $1.25 per unit

Production in 2021 = 30,600 units

Depreciation charge for 2021 = $1.25 x 30,600 = $38,250

Units-of-Output Depreciation charge for 2021 = $38,250

  1. Working hours –

Depreciation expense = depreciable rate x hours used

Depreciable rate = depreciable base/estimated hours

Depreciable base = cost – salvage value

Cost = $381,000

Salvage value = 18,000

Depreciable base = 381,000 – 18,000 = $363,000

Estimated hours = 25,000

Depreciable rate per hour = $363,000/25,000 = $14.52 per hour

Hours worked in 2021 = 2,650

Depreciation charge for 2021 = $14.52 x 2,650 = $38,478

Working hours depreciation charge for 2021 = $38,478

  1. Sum-of-the-years’-digits –

Depreciation expense = depreciable base x depreciation fraction

Depreciable base = cost – salvage value

Cost = $383,000

Salvage value = $18,000

Depreciable base = 381,000 – 18,000 = $363,000

Useful life = 10 years

Depreciation fraction = remaining useful life of machine/sum of the years

Remaining useful life = 9 years

Sum of the years = 10 + 9 + 8+7+6+5+4+3+2+1 = 55

Depreciation charge = 363,000 x 9/55 = $59,400

Sum-of-the-years’-digits Depreciation charge for 2021 = $59,400

  1. Declining balance –

Depreciation expense = cost x annual rate

First year depreciation expense = cost x annual rate

Cost = $381,000

Declining rate = 20%

2020 depreciation charge = 381,000 x 20% = $76,200

Depreciation for 2020 is for 8 months, May 1 – Dec 31, therefore depreciation charge = 76,200 x 8/12 = $50,800

Book value at Jan 1, 2021 = cost – accumulated depreciation

= 381,000 – 50,800 = $330,200

Depreciation charge for 2021 –

Book value x 20% = 330,200 x 20% = $66,040

Declining balance Depreciation charge for 2021 = $66,040